Circular No. 7384

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 2000R, 2000S, 2000T, 2000U, 2000V, 2000W, 2000X

A. Amadon and G. Blanc, Service de Physique des Particules (SPP), Centre d’Etude de Saclay (CES); J. Haissinski, Laboratoire de l’Accélérateur Linéaire, Orsay; and J. Rich, SPP/CES, on behalf of the EROS collaboration, report the discovery of seven supernovae discovered in the course of the automated supernovae search with the European Southern Observatory (ESO) 1-m Marly telescope at La Silla (limiting mag ∼ 21.5). Each supernova was found by subtraction of CCD images from reference images taken one year ago.

SN  2000 UT  α_{2000}  δ_{2000}  V  Offset
2000R Feb. 27 10^{h}01^{m}21.^{s}27 − 6^{o}33′57′′.4 19.4 1′′.6 W, 1′′.7 S
2000S Feb. 28 10 34 55.89 − 6 28 26.0 20.9 3′′.9 E, 10′′.9 S
2000T Mar. 3 12 36 20.35 −12 28 23.3 21.0 2′′.7 W, 0′′.9 S
2000U Mar. 4 12 45 27.90 −12 10 34.3 20.2 3′′.9 W, 1′′.5 N
2000V Mar. 6 12 03 28.64 − 5 47 15.7 20.6 3′′.8 W, 0′′.6 S
2000W Mar. 9 11 10 59.53 − 4 43 46.8 20.4 6′′.4 E, 3′′.5 S
2000X Mar. 10 11 30 45.41 − 5 12 16.8 20.2 8′′.0 E, 5′′.7 S

SN 2000R was confirmed via a CCD spectrum taken at the ESO 3.6-m telescope by C. Coutures (CES, ESO) on Mar. 6 that suggests it is a type-IIL supernova in its nebular phase with z = 0.01. SN 2000S in MCG –01-27-020 was also confirmed by Coutures with a CCD spectrum taken on Mar. 9, the data being consistent with a type-Ic event in its nebular phase, with z = 0.02. Spectra by Coutures on Mar. 6 and 9 of SNe 2000T and 2000V, respectively, are consistent with type-Ia supernovae near maxima and z = 0.18 and 0.10. SNe 2000U, 2000W, and 2000X were each confirmed by CCD images taken at the Marly telescope ∼ 2 days after discovery, and these three appear to have been found premaximum. SN 2000U increased in brightness by ∼ 0.5 mag between Mar. 4 and 12; SN 2000W increased by ∼ 0.6 mag between Mar. 8 and 15; SN 2000X increased by ∼ 0.5 mag between Mar. 10 and 15.

V382 VELORUM Visual magnitude estimates by A. Pearce, Nedlands, W. Australia:

Feb. 13.780 UT, 9.7; 16.823, 9.7; Mar. 5.769, 9.8; 6.790, 9.8; 7.781, 9.9; 13.771, 9.9; 15.781, 9.9; 17.794, 10.1.

2000 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green